Letty wants to retire at the end of 6 years and thinks that that retirement party is going to cost $9375. If the interest rate is 10%, how much should Letty save each year?
Answer #1

Desired Sum in 6 years = $9,375
Interest Rate = 10%
Period = 6 years

Let Annual Deposit be $x

$x*1.10^5 + $x*1.10^4 + $x*1.10^3 + $x*1.10^2 + $x*1.10 + $x = $9,375
$x * (1.10^6 - 1) / 0.10 = $9,375
$x * 7.71561 = $9,375
$x = $1,215.07

So, annual deposit is $1,215.07